Alternative Exercise 4.61:

Constants Two forces have the same magnitude F

What is the angle between the two vectors if their sum has a magnitude of 2F? Request Answer Part B tween the two vectors if theit sum has a magnitude of v2F9

Part C What is the angle between the two vectors if their sum has a magnitude of zero? 4% Submit Request Answer

Need Part A, Part B, and Part C
Answer #1

Let A be the vector having magnitude F

B be the vector having magnitude F

Now by the vector law

(A+B)^2=A^2+B^2+2*A*B*cosTheta ................(1)

Now in part b

Sum of both vectors is 2F

A+B=2F

(1)->  (2F)-=F2 + F2 + 2 * F * F * coso

4=2+2cosTheta

2=1+cosTheta

cosTheta =1

Theta =0

Hence the Angle between two vector is 0

For part c

Now sum of both vectors is 0

A+B=0

(A+B)^2=A^2+B^2+2*A*B*cosTheta

0=F2 + F2 +2 * F * F * co:0

0=2+2cosTheta

0-1 cose

cose- -1

Theta =180°

Is answer